According to the USA Swimming association, the distribution of finishing time, in seconds, for a 50-yard Freestyle competition for girls 8 years and under is approximately bell-shaped with a mean of 41 seconds and a standard deviation of 5 seconds.

1. Natalie is a 7 year old girl that participated in a 50-yard Freestyle competition and her finishing time is 36 seconds. Based on the stated model, approximately, what percentage of the 50-yard Freestyle finishing times for girls 8 years and under are expected to be slower than Natalie?

  • 95%

  • 68%

  • 32%

  • 16%

  • 84%

  • 2.5%

  • 5%

2. Find the value to complete the following statement:
The fastest 2.5% of female swimmers 8 years and under will complete a 50-yard Freestyle competition in at most _________ seconds, approximately.

  • 26

  • 31

  • 36

  • 41

  • 46

  • 51

  • 56
Answer #1

We have given that

μ = 41, σ 5

X: finishing time

1.  Natalie's finishing time is 36 seconds.

Now we have to calculate

X-μ / 36-41 P(X < 36) = P( P(Z <-1) 0.1587 ะ 0.16 ANS : 16%

2. The fastest 2.5% of female swimmers 8 years and under will complete a 50-yard Freestyle competition in at most 31 seconds, approximately.

Since we have to calculate X using fastest 2.5% i.e. 0.025

i.e. F(z) = PlZ < ) = 0,025 = F-1 (0.025) =-1.96

So ~X=mu+z*sigma=41+(-1.96)*5=31.2002approx 31